OpenStudy (anonymous):

Find the coordinates of the missing endpoint given that B is the midpoint of AC

jimthompson5910 (jim_thompson5910):

This is the full problem?

OpenStudy (anonymous):

What do u mean?

jimthompson5910 (jim_thompson5910):

Does it give you the coordinates of A, B or C?

OpenStudy (anonymous):

The coordinates are C(0,0), B(5,-6)

jimthompson5910 (jim_thompson5910):

So let A be (x,y)

jimthompson5910 (jim_thompson5910):

this means that (x+0)/2 = 5 and (y+0)/2 = -6

OpenStudy (anonymous):

So it would be (5,-6) for the answer

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

(x+0)/2 = 5 (x)/2 = 5 x/2 = 5 x = 2*5 x = 10 do the same for (y+0)/2 = -6 to find y

OpenStudy (anonymous):

Ok